Suppose that the index model for stocks A and B is estimated from excess returns with the following results:

RA = 3.00% + 1.05RM + eA

RB = -1.20% + 1.20RM + eB

σM = 29%; R-squareA = 0.29; R-squareB = 0.14

Assume you create portfolio P with investment proportions of 0.60 in A and 0.40 in B.a. What is the standard deviation of the portfolio? (Do not round your intermediate calculations. Round your answer to 2 deci
Answer #1

Given information:

RA = 3.00% + 1.05RM + eA This implies \beta_A = 1.05

RB = -1.20% + 1.20RM + eB This implies \beta_B = 1.20

σM = 29%; R-squareA = 0.29; R-squareB = 0.14

a.What is the standard deviation of the portfolio?

To calculate standard deviation of portfolio, first we need to determine standard deviation of A and B.

The standard deviation can be derived from R2 equation

R_A^2 = Explained variance / Total variance = \beta_A^2 * \sigma_M^2 / \sigma_A^2

0.29 = 1.052 * 0.292 / \sigma_A^2

solving we get \sigma_A = 0.5654

Similarly for stock B,

R_B^2 = Explained variance / Total variance = \beta_B^2 * \sigma_M^2 / \sigma_B^2

0.14 = 1.22 *0.292 / \sigma_B^2

solving we get \sigma_B = 0.9301

Covariance between A and B is given by the formula:

Cov (RA, RB) = \beta_A * \beta_B * \sigma_M^2 = 1.05 * 1.2 * 0.292 = 0.106

Variance of portfolio = w* 0+ W *OB+2 * WA* WB* Cou(A,B)

= 0.62*0.56542+0.42*0.93012+2*0.6*0.4*0.106

= 0.3044

Hence standard deviation of portfolio = square root (0.3044) = 0.5517

b. Calculate beta of portfoli0

R(Portfolio) = 0.6RA + 0.4RB

=0.6*(3+1.05RM +eA) +0.4 *(-1.20 + 1.20RM + eB)

= 0.18 + 0.630RM +0.6eA - 0.48 + 0.48RM + 0.4eB

= 0.66+1.11 RM + 0.6eA + 0.4eB

The co-efficient of RM is the beta of the portfolio. Hence beta of portfolio is \beta_P = 1.11

c. What is the firm specific variance of the portfolio?

Firm specific risk = Total risk - Systematic risk

We have already calculate total risk of portfolio in part a. Now we need to calculate systematic risk.

Systematic risk of portfolio = BOM

Beta of portfolio = 1.11 which we calculated in part b.

Hence systematic risk of portfolio = 1.112*0.292 = 0.1036

substitute this value in the formula above:

Firm specific risk = 0.55172 - 0.1036 = 0.2001

Part d:

What is the covariance between portfolio and market index?

Cov (RP, RM) = Bp * o

substitute the values

Cov (RP, RM) = 1.11 * 0.292 = 0.09335

Answer:

Covariance between portfolio and market is 0.09335